why co+ has bond order of 3.5?

Grade:12th Pass

1 Answers

ruchi yadav
10 years ago
The BO in CO+ is 3.5 because the order of the orbitals is not the same as it would be for a homonuclear diatomic molecule such as O2where both atoms are the same. In the case of CO, the 2s atomic orbital on Oxygen is much lower in energy than the 2s atomic orbital in carbon. The discrepancy in energies allows the 2p-pi and 2p-sigma molecular orbitals to sink lower than the 2s-sigma* orbital in the MO diagram. For this case the calculations would be:

For CO:

2s-sigma22p-pix22p-piy22p-sigmaz22s-sigma*2

BO = (8-2)/2 = 3

For CO+:

2s-sigma22p-pix22p-piy22p-sigmaz22s-sigma*1

BO = (8-1)/2 = 3.5


Thank You
Ruchi
Askiitians Faculty